### v1

# Study of charged Lepton Flavor Violation in electron muon interactions

Abstract
With the improvement of muon acceleration technology, it has received great interest to exploit high-energy muon beams for collision or target experiments. We investigate possible charged Lepton Flavor Violation (cLFV) processes mediated by an extra massive neutral gauge boson $Z^{\prime}$ in electron muon interactions, either at a proposed electron muon collider or in a fixed target experiment with high-energy muon beam hitting electrons in the target. Based on Monte Carlo calculations and fast detector simulations, we study in detail our signal and possible backgrounds, giving the sensitivity results of cLFV signals at the 90% confidence level. Compared with current and prospective limits set by other experiments, electron muon interactions demonstrate significant advantages in the cLFV coupling strength sensitivity with $\tau$ in the final states. In addition, a special cLFV coupling combination, $\lambda_{e\mu}\times\lambda_{\mu\mu}$, can also be probed in our proposal.
Keywords:
electron muon interaction, charged Lepton Flavor Violation, extra gauge boson
arxiv: 1234.56789
1 Introduction
Collider experiment serves as a crucial tool for precision measurement of the standard model (SM) and search for new physics beyond the SM (BSM), with the experimental technology constantly being improved and evolved. In the near future, the High-Luminosity Large Hadron Collider (HL-LHC) Apollinari:2017lan , the Future Circular Collider (FCC) FCC:2018byv ; FCC:2018evy ; FCC:2018vvp or the Circular Electron Positron Collider (CEPC) CEPCStudyGroup:2018rmc ; CEPCStudyGroup:2018ghi may become important instruments for the next generation high energy frontier research. While recently with the continuous development of muon acceleration technology, the muon collider has also become an increasingly popular consideration. Since it integrates the advantages of electron colliders and hadron colliders, the muon collider may become a golden factory for studying various new physics processes Aime:2022flm .
On the other hand, high energy muon beams can also be used to create electron muon collisions. As early as ten to thirty years ago, numerous research efforts have already focused on the potential of electron muon collider Barger:1997dv ; Choi:1997bm ; Montero:1998sv ; Cvetic:1999fk ; Almeida:2000qd ; Singhal:2007hw . In recent years, as the construction of high energy lepton collider gradually became feasible from the engineering standpoint, interest in electron-muon collision has been reignited Bouzas:2023vba ; Bossi:2020yne ; Lu:2020dkx ; Bouzas:2021sif .
Recently, a new collider proposal, $\mu$TRISTAN, has been proposed based on an ultra-cold muon technology developed for the muon $(g-2)$ experiment at J-PARC Hamada:2022mua . It includes a $\mu^{+}\mu^{+}$ collider and a $\mu^{+}e^{-}$ collider, in which we are interested in the later. The main parameters from the $\mu$TRISTAN $e\mu$ collider proposal Hamada:2022mua ; Akturk:2024jbl are listed in Tab.1. According to several phenomenological study, $\mu$TRISTAN may have certain potentials on measurements related to Higgs and new physics searches Das:2024gfg ; Lichtenstein:2023iut ; Hamada:2022uyn .
Table 1: The main parameters of the $\mu$TRISTAN $e\mu$ collider.
Parameter Electron Anti-muon
Beam energy 30 GeV 1 TeV
Polarization 70% ¿25%
Particles per bunch ($10^{10}$) 6.2 1.4
Luminosity 4.6×1033 cm-1s-1
Collision frequency 4×106 Hz
Meanwhile, utilizing high-density muon beams to strike fixed targets can also provide a possibility to search for new physics. Many such attempts have been proposed, including the Muon Missing Momentum ($M^{3}$) proposal at Fermilab kahn2018m3 , and a recent idea of searching for muonic force carriers by using ATLAS detector as a fixed target galon2020searching .
In this study we also investigate a fixed electron-target experiment with a muon beam in addition to the $e\mu$ collider. Since lepton flavor in the initial state is non-zero, electron muon interaction can strongly avoid many potential background processes which would occur at different-sign muon colliders or electron-positron colliders, thus possessing higher sensitivity to new physics signals, typically the charged Lepton Flavor Violation (cLFV) processes.
In the SM framework, the cLFV processes are strongly suppressed due to the tiny mass of neutrinos, hence unobservable in the current experiments yet. However, it may be much enhanced in various BSM models, such as super-symmetry (SUSY) Barbier:2004ez , leptoquark Dorsner:2016wpm , two-Higgs-doublet Branco:2011iw , and the heavy neutral gauge bosons $Z^{\prime}$ Langacker:2008yv studied in this paper. In the past decades, searches for the cLFV process were performed in different channels with several approaches, typically the high intensity muon-based experiments including $\mu^{+}\to e^{+}\gamma$ (MEG) MEG:2016leq , $\mu^{+}\to e^{+}e^{+}e^{-}$ (SINDRUM) SINDRUM:1987nra and $\mu^{-}N\to e^{-}N$ (SINDRUM \@slowromancapii@) SINDRUMII:2006dvw ; SINDRUMII:1993gxf ; SINDRUMII:1996fti ; SINDRUMII:1998mwd , as well as the collider-based searches for cLFV decays of $Z$ ATLAS:2014vur ; OPAL:1995grn ; DELPHI:1996iox , Higgs ATLAS:2019pmk ; CMS:2017con and several hadron resonances KTeV:2007cvy ; CLEO:2008lxu ; BaBar:2021loj ; BESIII:2022exh . Meanwhile, there will be continuous new experiments conducted in the near future to constantly improve the existing limits, such as MEG\@slowromancapii@ MEGII:2018kmf , Mu3e Mu3e:2020gyw , COMET COMET:2018auw and Mu2e Mu2e:2014fns .
In this study, we consider the cLFV processes based on the interactions of electron and muon in two scenarios: asymmetric electron muon collision at the $e\mu$ collider and fixed electron-target experiment striking with the muon beam. For the former case, the center-of-mass energy includes the energy point of $\mu$TRISTAN and even higher. While for the latter case, we investigate the muon energy around several tens of GeV to test the $Z^{\prime}$ couplings at the low energy bound.
2 Physics processes and Monte Carlo simulation
2.1 cLFV in $Z^{\prime}$ model
By introducing an additional $U(1)$ gauge symmetry into the SM framework, it will correspond to a neutral gauge boson $Z^{\prime}$. Since the heavy neutral gauge bosons are predicted in many BSM models, it may be one of the most motivated extensions of the SM Langacker:2008yv ; Langacker:2000ju ; Buras:2021btx .
In this study, $Z^{\prime}$ is considered to have the same coupling and chiral structure as the standard model $Z^{0}$, but allows for the lepton flavor violation, similarly as Ref. ATLAS:2018mrn ; Li:2023lin . The coupling strength of the $Z^{\prime}$ and leptons can be described by a matrix $\lambda$ as Eq. 1
$\lambda=\left(\begin{array}[]{lll}\lambda_{ee}&\lambda_{e\mu}&\lambda_{e\tau}\\ \lambda_{\mu e}&\lambda_{\mu\mu}&\lambda_{\mu\tau}\\ \lambda_{\tau e}&\lambda_{\tau\mu}&\lambda_{\tau\tau}\end{array}\right).$ (1)
Generally, it represents the strength of the cLFV couplings relative to the SM couplings, assuming that the diagonal elements are 1, while the off-diagonal elements are usually a higher order of magnitude. Therefore, the cases that break lepton flavor conservation twice would not be considered in the further study. After introducing the $Z^{\prime}$ boson, the cLFV processes mentioned in Sec.1 can be enhanced by the diagrams shown in Fig.1. The branching ratio limits would be transformed to the coupling $\lambda_{ij}$ Bernabeu:1993ta ; Langacker:2000ju and compared with our results based on $e\mu$ interaction.
Figure 1: The Feynman diagrams of several $Z^{\prime}$ mediated cLFV processes.
2.2 Physics processes
The signal cLFV processes studied in this paper are listed in Tab.2. There are two diagrams for the $Z^{\prime}$ mediated cLFV process $\mu^{+}e^{-}\to l^{+}l^{-}$, as shown in Fig.2 (taking $\mu^{+}e^{-}\to e^{+}e^{-}$ as an example). In particular, the s-channel is not included in the processes $\mu^{+}e^{-}\to\mu^{+}\tau^{-}$. In the simulation, all coupling strengths are considered as 1. And the mass of $Z^{\prime}$ floats from 0.2 GeV to 5 TeV in the electron muon collision experiment, and within 0.50 GeV in the electron-target experiment with a muon beam.
Figure 2: The Feynman diagrams of the process $\mu^{+}e^{-}\to e^{+}e^{-}$: (a) s-channel and (b) t-channel.
While several background processes may occur on the collider and affect the signal that we are interested in, these processes include the standard model backgrounds divided by the number of final state particles and accidental background caused by particle mis-identification. In conclusion, the specific signals and their background processes are shown in Tab.2.
Table 2: Specific processes of signal and SM background
Signal Process SM and accidental backgrounds.
$\mu^{+}e^{-}\to e^{+}e^{-}\nu_{e}\bar{\nu}_{\mu}$
  $\mu^{+}e^{-}\to e^{+}e^{-}$ $\mu^{+}e^{-}\to e^{+}e^{-}\nu_{e}\bar{\nu}_{\mu}\nu\bar{\nu}$
$\mu^{+}e^{-}\to\mu^{+}e^{-}$
$\mu^{+}e^{-}\to\mu^{+}\mu^{-}\nu_{e}\bar{\nu}_{\mu}$
  $\mu^{+}e^{-}\to\mu^{+}\mu^{-}$ $\mu^{+}e^{-}\to\mu^{+}\mu^{-}\nu_{e}\bar{\nu}_{\mu}\nu\bar{\nu}$
$\mu^{+}e^{-}\to\mu^{+}\mu^{-}$
  $\mu^{+}e^{-}\to\mu^{+}\tau^{-}$ $\mu^{+}e^{-}\to\mu^{+}\tau^{-}\nu_{e}\bar{\nu}_{\mu}$
$\mu^{+}e^{-}\to\mu^{+}\tau^{-}\nu_{e}\bar{\nu}_{\mu}\nu\bar{\nu}$
2.3 Event generation and simulation
Both signal and background events are simulated by MadGraph5_aMC@NLO(MG) Madgraph version 3.1.1, which is one of the key tools for Monte Carlo event generation in high energy physics, then showered and hadronized by Pythia8 Pythia . Next, Delphes Delphes version 3.5.1 is utilized to simulate detector effects with the default configuration card for the detector at the muon collider.
In the Monte Carlo generation, some preliminary requirements are applied to remove the physically unreasonable events. In $e\mu$ collisions, the transverse momentum of charged leptons is required to satisfy $p_{T}>$ 10 GeV and the absolute pseudo-rapidity of charged leptons $\lvert\eta\rvert>$ 2.5. While in the muon-beam electron-target simulation, the filtering criteria of $p_{T}$ and $\lvert\eta\rvert$ would be relaxed. Then in detector simulation, parameters such as efficiency of particle detection are set according to the Delphes cards.
There is some difference between the $\tau$ simulation and $e/\mu$. For those final states with $\tau$ in Tab.2, although they can go through any decay chains in reality, in this study we only consider the hadronic decay channels (about 60% of the total decay) and reconstruct it by the Jet information.
3 Statistical analysis and sensitivity result
3.1 Asymmetric collision
In this scenario we consider two kinds of asymmetric collision with electron and anti-muon beam: $E_{e}=30$ GeV and $E_{\mu}=1$ TeV ($\sqrt{s}=346$ GeV), or $E_{e}=200$ GeV and $E_{\mu}=3$ TeV ($\sqrt{s}=1.55$ TeV). The former is based on the proposal of $\mu$TRISTAN (the polarization of each beam is not considered), and the latter is a higher energy assumption according to other current beam designs.
3.1.1 Background study
After setting the preliminary requirements as mentioned in Sec.2.3, signal candidate events should be with the same charged leptons corresponding to the signal in the final state. Since the initial flavor in $e\mu$ collision is non-zero, vast majority of the SM background are forbidden, while the remaining portion also exhibits significant kinematic differences from the signal processes, especially in $ee$ or $\mu\mu$ channel. The invariant mass distributions of the final state dileptons of these two channels are shown in Fig.LABEL:fig:mlldistribution1.
Only considering the interval near the center-of-mass energy, the SM background values are extremely low. And compared with a similar study on the different-signs election or muon collider Li:2023lin , $e\mu$ collision has a cleaner signal window. Due to the low level of the SM background, we also investigated the accidental background caused by $e\mu$ mis-identification, where a final state muon is assigned the mass of the electron, or vice versa. Typically it would let the $e\mu$ scattering process coming into the background of $\mu^{+}e^{-}\to e^{+}e^{-}$ and $\mu^{+}e^{-}\to\mu^{+}\mu^{-}$. The probability of $e\mu$ mis-identification is set as 10-6. The invariant mass distribution of this process is extremely close to the signal.
while for $\mu^{+}e^{-}\to\mu^{+}\tau^{-}$ since the reconstruction of $\tau$ would inevitably result in a certain loss of energy, there is a considerable overlap in the signal and background distributions. It will be optimized in the next section.
3.1.2 Sensitivity result
Based on the distributions of the signal and background, we truncate the invariant mass to remove the events with significantly deviating from center-of-mass energy. The specific truncation point in $\tau$ channels will be determined by scanning and selecting the maximum value of $S/(a/2+\sqrt{B})$, where $S$ is the number of signal processes, $B$ is the total number of the weighted background and $a$ is the significant value which is considered as 3 Punzi:2003bu . The weight is defined by $n_{x}=\sigma_{x}L/N$, where $\sigma_{x}$ is the cross section of each process, $L$ is the luminosity and $N$ is the generated number.
Then the binned histograms of leptons $p_{T}$ distributions are utilized for the statistic analysis. The test statistics $Z_{i}$ is calculated by $Z_{i}:=2[n_{i}-b_{i}+b_{i}\ln{(n_{i}/b_{i})}]$ for 90% confidence level (C.L.) exclusion, where $i$ is the index of each bin, $n$ is the weighted number of observed events including signal and background, and $b$ is the weighted number of background Cowan:2010js . Then the total $Z=\Sigma_{i}Z_{i}$ would be subject to a $\chi^{2}$ distribution. The degree of freedom is defined by the number of bins. By iteration, we can obtain the signal cross section of 90% C.L. exclusion in Fig.4.
Then based on MG calculation we can get the corresponding value of $\lambda_{ll^{\prime}}\times\lambda_{\rm SM}$, as shown in Fig.5. The $\lambda_{e\mu}$ and $\lambda_{e\tau}$ results are calculated at two energy points, and the luminosity is considered as 0.3 ab-1 or 5 ab-1. Several current limits and prospective limits are also included to compare with our results. The constraint of branching ratio from other experiments are concluded in Tab.3. It is sensible that the limits given by other experiments in $\tau$ channels are much more conservative than in $e\mu$ channels. But in our collider study, they may be similar since the main influencing factor here is the signal cross section while the cross sections of different signals are quite similar.
Figure 4: Cross section of each cLFV signal process of 90% C.L. exclusion relative to the luminosity of $e\mu$ collider.
Table 3: Summary of current and prospective limits from other experiments at 90% C.L.
Coupling Channel Constraint of branching ratio
Current prospective
$\mu N\to eN$ $6.1\times 10^{-13}$ SINDRUMII:1998mwd $3.0\times 10^{-17}$ COMET:2018auw
$\lambda_{e\mu}$ $\mu\to eee$ $1.0\times 10^{-13}$ SINDRUM:1987nra $1.0\times 10^{-16}$ Mu3e:2020gyw
$\mu\to e\gamma$ $4.2\times 10^{-13}$ MEG:2016leq $6.0\times 10^{-14}$ MEGII:2018kmf
$\tau\to e\gamma$ $3.3\times 10^{-8}$ Hayasaka:2010np $9.0\times 10^{-9}$ Banerjee:2022vdd
$\lambda_{e\tau}$ $\tau\to\mu\mu e$ $2.7\times 10^{-8}$ Hayasaka:2010np $4.5\times 10^{-10}$ Banerjee:2022vdd
$\tau\to eee$ $2.7\times 10^{-8}$ Hayasaka:2010np $4.7\times 10^{-10}$ Banerjee:2022vdd
Figure 5: The 90% C.L. upper limit of $\lambda_{e\mu}\times\lambda_{\rm SM}$ (a) and $\lambda_{e\tau}\times\lambda_{\rm SM}$ (b). The curves are graphed with respect to $M_{Z^{{}^{\prime}}}$, representing the limits of the cross section times branching ratio. Additionally, exclusion lines from both present low-energy experiments (shown as solid lines) and future experiments (shown as dashed lines) are included in the plot.
In the $\lambda_{e\mu}$ channel, the strictest constraint comes from the results of $\mu-e$ conversion, and our results have weaker competitiveness among those high intensity muon-based experiments. While in the $\lambda_{e\tau}$ channel, current limits of $\tau\to eee$ and $\tau\to e\mu\mu$ perform the best across the entire interval, but the orders of magnitude are much lower than the results of $\lambda_{e\mu}$. Compared with our results, the constraints of 1.55 TeV $e\mu$ collider are more stringent than other existing results when $M_{Z^{\prime}}>0.5$ TeV, and of 346 GeV collider it is about the entire $M_{Z^{\prime}}$ interval. Even compared with the prospective limits of those processes on Belle II, our results still have certain advantages around the resonance region.
On the other side, comparing the results of $e\mu$ collision and $e^{+}e^{-}$ or $\mu^{+}\mu^{-}$ collisions Li:2023lin , based on the same luminosity, the results of $e\mu$ are better than $\mu^{+}\mu^{-}$ by an order of magnitude at their respective resonance points. But it cannot be ignored that the center-of-mass energy of $e\mu$ is relatively low, and the results of other experiments are more advantageous in the low-energy region. Not only that, for $e\mu$ collider it is difficult to let the luminosity reach the same level as $\mu^{+}\mu^{-}$ collider in reality. Therefore, the advantages of these two collision scenarios still need to be considered based on more practical factors.
To be precise that what we are comparing in Fig.5(a) and Fig.5(b) is the coupling $\lambda_{e\mu}(\lambda_{e\tau})\times\lambda_{\rm SM}$. Then based on the assumption that $Z^{\prime}$ has the same coupling structures and strengths as the standard model $Z^{0}$ as mentioned in Sec.2.1, that is, $\lambda_{\rm SM}=1$, we can naturally obtain the estimation of $\lambda_{e\mu}$ or $\lambda_{e\tau}$. But more strictly speaking, this requirement may not exist in more universal models, and we need to consider the biases of standard model couplings. In this way, each process may give a product of different coupling strength, for example $\lambda_{e\mu}\times\lambda_{ee}$ ($\mu\to eee$), and $\lambda_{e\mu}\times\lambda_{ll}$ ($\mu\to e\gamma$) where $l$ represent $e$ or $\mu$. While for the high intensity muon-based experiments, they were unable to measure the coupling $\lambda_{e\mu}\times\lambda_{\mu\mu}$. This coupling requires the existence of a $Z^{\prime}-e-\mu$ vertex and a $Z^{\prime}-\mu-\mu$ vertex, which is almost impossible to achieve except for the direct interactions of leptons. On the $e\mu$ collider, we can obtain the measurement of this coupling through $\mu^{+}e^{-}\to\mu^{+}\mu^{-}$, as shown in Fig.6, indicating that it has similar sensitivity as $\mu^{+}e^{-}\to e^{+}e^{-}$.
Figure 6: Sensitivity of the coupling strength $\lambda_{\mu\mu}\times\lambda_{e\mu}$ from $\mu^{+}e^{-}\to\mu^{+}\mu^{-}$ channel.
3.2 Electron-target experiment with a muon beam
Now we will focus on the low-energy region, where we conducte Monte Carlo simulations to investigate muon-target processes involving $\mu^{+}e^{-}\to e^{+}e^{-}$ and $\mu^{+}e^{-}\to\mu^{+}\mu^{-}$. It is essential to highlight that the masses of muon and electron cannot be disregarded, imposing a lower limit on the incident muon’s beam energy, equivalent to its mass. Additionally, the target energy $E_{cm}=\sqrt{2E_{\mu}m_{e}+{m_{\mu}}^{2}+{m_{e}}^{2}}$ possesses its own lower-energy threshold. Consequently, in our simulations, we vary the $M_{Z^{{}^{\prime}}}$ in three distinct sets for each process, ensuring that the muon energy is scanned as close to the lower energy limit as feasible. The outcomes are shown in Fig.LABEL:fig:resonance and Tab.4. Remarkably, a pronounced resonance in the target energy is observed in proximity to $M_{Z^{{}^{\prime}}}$.
Process $M_{Z^{{}^{\prime}}}$ / $\mathrm{GeV}$ $E_{\mu}$ / $\mathrm{GeV}$ $E_{e}$ / $\mathrm{MeV}$ $E_{cm}$ / $\mathrm{GeV}$
$\mu^{+}e^{-}\to e^{+}e^{-}$ 0.11 0.93 0.511 0.1101
0.15 11.1 0.511 0.1501
0.20 28.2 0.511 0.1996
$\mu^{+}e^{-}\to\mu^{+}\mu^{-}$ 0.22 33.6 0.511 0.2200
0.25 50.2 0.511 0.2499
0.30 77.2 0.511 0.2998
Table 4: Resonant collision energy of process $\mu^{+}e^{-}\to e^{+}e^{-}$ and $\mu^{+}e^{-}\to\mu^{+}\mu^{-}$ with different $M_{Z^{{}^{\prime}}}$.
Considering the processes we are studying, taking $\mu^{+}e^{-}\to\mu^{+}\mu^{-}$ as an example, the main background process corresponding here is $\mu^{+}+e^{-}\to\mu^{+}+\mu^{-}+\gamma$ kahn2018m3 . Correspondingly, its relative rate is $10^{-8}$, therefore we can conduct background-free experimental estimation. Due to $\sigma\propto(\lambda_{e\mu}\times\lambda_{ll})^{2}$and event rate $R=L\cdot\sigma=\frac{dN}{dt}\cdot n_{2}\cdot dx\cdot\sigma$, coupling limits estimates can be made based on the reaction cross-section. As a rough estimate, assuming a 10 cm thick lead target, the incidence rate is about $\frac{dN}{dt}\sim 10^{6}$, the electron number density of lead is about $n_{2}\sim 10^{24}$, and $1\mathrm{y}\sim 10^{7}$\mathrm{s}$$. From this it can be obtained 90% C.L. exclusion lines on the couplings $\lambda_{ee}$ and $\lambda_{\mu\mu}$ products the diagonal coupling $\lambda_{e\mu}$.
The current and prospective limits from low-energy experiments are converted to the coupling limits on $\lambda_{e\mu}\times\lambda_{SM}$ to compare with our results. As shown in Fig.LABEL:fig:3.2.1, in comparison to current results and prospect MEG\@slowromancapii@ experiment, our $ee$ channel simulation can give more stringent upper limits of coupling $\lambda_{\mu e}$. Furthermore, as mentioned in Sec.3.1.2. we also can set new limits on coupling $\lambda_{\mu\mu}\times\lambda_{\mu e}$ as shown in Fig.LABEL:fig:3.2.2, which have never been obtained from experiments yet.
4 Conclusion
With the continuous development of muon technology, in addition to building high-energy muon collider, there is also certain research prospect for the $e\mu$ interaction. In this work, we investigated the cLFV process propagated by massive neutral gauge bosons ($Z^{\prime}$) in $e\mu$ collision and electron-target experiment with a muon beam, in order to explore the potential of $e\mu$ interactions in new physics searches. We conduct the simulation studies on the cLFV processes $\mu^{+}e^{-}\to e^{+}e^{-}$, $\mu^{+}e^{-}\to\mu^{+}\mu^{-}$ and $\mu^{+}e^{-}\to\mu^{+}\tau^{-}$, using MadGraph5_aMC@NLO, Pythia8 and Delphes. Then we provide the coupling strength $\lambda_{e\mu}$ and $\lambda_{e\tau}$ at 90% C.L. of different $M_{Z^{\prime}}$. By comparing the sensitivity results with current and prospective limits, it is shown that $e\mu$ interactions have certain research advantages in $\tau$ channel for the heavy $Z^{\prime}$. Furthermore, through the direct interaction of leptons, a unique $\lambda_{e\mu}\times\lambda_{\mu}$ process can be measured, which is unmatched by most other experiments.
Acknowledgements.
This work is supported in part by the National Natural Science Foundation of China under Grant No. 12150005, 12075004, 12175321, 12061141003; the National Key Research and Development Program of China under Grant No. 2018YFA0403900; National College Students Innovation and Entrepreneurship Training Program, Sun Yat-sen University; State Key Laboratory of Nuclear Physics and Technology, Peking University under Grant No. NPT2020KFY04, NPT2020KFY05.